### v1

# Large Language Models (LLMs) for Wireless Networks: An Overview from the Prompt Engineering Perspective

Abstract
Large language models (LLMs) have been successfully applied to many fields, showing outstanding comprehension and reasoning capabilities. Despite their great potential, LLMs usually require dedicated fine-tuning for domain-specific applications such as wireless networks. Such domain adaptation can be extremely demanding for computational resources, while many network devices have limited computation capacity. To this end, this work explores LLM-enabled wireless networks from the prompt engineering perspective, i.e., designing input prompts to guide LLMs to generate desired output. Compared with other approaches to using LLMs, prompt engineering is more flexible and resource-efficient, which better aligns with wireless network features. In particular, we first introduce LLM fundamentals and compare different prompting techniques such as in-context learning, chain-of-thought, and self-refinement. Then we propose two novel prompting schemes for network applications, namely iterative prompting for wireless network optimization, and self-refined prompting for network prediction. The case study shows that the proposed schemes can achieve comparable performance as conventional machine learning techniques, and our prompting-based methods avoid the complexity of dedicated model training and fine-tuning, which is one of the key bottlenecks of existing machine learning techniques.
Index Terms:
Large language models, wireless networks, prompt engineering
I Introduction
TABLE I: Summary of various LLM usage approaches for wireless networks.
Approaches & Steps Main features Advantages Potential difficulties Computational resources & Time costs Possible wireless network applications
Pre-training LLMs Training LLMs from scratch on large-scale datasets, typically involving hundreds of billions of tokens for the next token prediction. It is a foundational step of LLM development and usage, including dataset collection and model training. Pre-training enables LLMs fundamental and crucial capabilities, such as comprehension, reasoning, zero-shot generation, and instruction following, which can address various downstream tasks. Pre-training is extremely resource-intensive, requiring massive labelled data, long training times, substantial energy consumption, and a tremendous amount of hardware accelerators like GPUs or TPUs. High Thermal Design Power GPUs such NVIDIA H100100100100-$80$G or A100100100100-$80$G are required. For example, training Llama3.13.13.13.1-$405$B took approximately $30.84$M GPU hours with $15$T tokens from publicly available sources and $25$M synthetic data. A network-specific LLM indicates great benefits, but training a network-specific LLM from scratch may be inappropriate due to the computational resource requirements.
Fine-tuning Fine-tuning refers to adapting a pre-trained LLM to specific tasks by partially or completely updating the model weights or introducing additional adapter modules to fit smaller and domain-specific datasets. Fine-tuning significantly reduces time and computational cost compared to pre-training, producing LLM models tailored to specific tasks. Fine-tuning has to consider the risk of overfitting and losing parts of the generalizability. Labelled data or pairwise comparison data may be needed, involving additional human labour force. The cost may vary between model sizes and fine-tuning methods. Typically, if full fine-tuning a 7B parameter model with 8 A100100100100-$80$G GPUs, it could be $100$-$200$ GPU hours on a $100$M token dataset (assuming a maximum length of $512$ tokens). Fine-tuning is a realistic approach to apply LLMs to wireless networks, adapting a general-domain LLM to specific tasks such as network troubleshooting, project coding, and network configuration.
Retrieval augmented generation (RAG) RAG combines LLMs with an external knowledge base, enabling the model to retrieve relevant information during inference. It enhances LLM’s capability to provide accurate and up-to-date responses, particularly in knowledge-intensive tasks. RAG can improve LLM’s response accuracy by leveraging external knowledge sources. It is particularly useful when the model’s training data is outdated or lacks specific information. RAG involves extra integration complexity, as it requires efficient retrieval algorithms since the inference latency may increase if the retrieval process is not optimized. Moreover, RAG highly depends on a well-structured knowledge base, which cannot be easily created and maintained. The computational cost depends on the retrieval method and the size of the knowledge base. Inference may take seconds to minutes per query, depending on the size of the base language model. Requires a balance between retrieval speed and accuracy, which can influence the choice of hardware, e.g., using faster storage for the knowledge base. Considering the cost of fine-tuning, RAG is a more practical approach for implementing LLMs, especially when existing network-specific datasets are available. It ensures the LLM has access to the most relevant and up-to-date network data, improving its accuracy for addressing the latest network tasks.
Prompt engineering Prompting engineering refers to crafting specific input prompts to guide a pre-trained LLM to generate desired outputs. It leverages the existing capabilities of the LLM without any model retraining or tuning. Prompt engineering has low computational costs and quick implementation. It is also flexible and can be easily adapted to different tasks without additional training. It requires expertise in crafting effective prompts, as the quality of the output heavily depends on how the prompt is formulated. It is also limited by the inherent capabilities of the pre-trained LLMs. It has minimal computational costs and often requires just a single GPU or a few GPUs. Crafting and experimenting with different prompts may take only a few hours. Time costs are significantly lower compared to other methods, making it suitable for rapid prototyping and testing. Prompt engineering enables rapid decision-making and responses in dynamic wireless network environments. It aligns well with many network tasks, especially considering the limited computational resources of network devices.
  • 1
    Note that these approaches can be combined to achieve better performance, e.g., fine-tuning network LLMs and then prompting.
As a sub-field of generative AI, large language models (LLMs) have received considerable interest from industry and academia [1]. After being trained on a large number of datasets, LLMs have shown revolutionary comprehension and reasoning capabilities. The advancement of generative AI and LLMs also provides promising opportunities for 6G wireless networks, including integrated satellite-aerial-terrestrial networks [2], vehicular networks [3], semantic communication [4], etc. Despite the great potential, integrating LLMs into wireless networks still faces several challenges. Firstly, wireless networks are complex large-scale systems with various knowledge domains, i.e., signal processing and transmission, network architecture and design, protocol, standards, and so on. Applying general-domain LLMs directly to domain-specific network tasks may lead to poor performance. Secondly, LLMs are extremely demanding for computational resources. LLM pre-training and fine-tuning are usually implemented on high-performance GPUs such as NVIDIA A100 and H100. However, wireless network devices usually have limited computational and storage capacities, which may prevent practical LLM network applications. In addition, LLMs can be applied in various approaches, including pre-training models from scratch, fine-tuning LLMs for domain-specific tasks, retrieval augmented generation (RAG), prompt engineering, etc. Each method has unique features and potential difficulties, and it is critical to identify the most efficient methods to adapt LLMs to wireless networks.
Given the above opportunities and challenges, this work focuses on prompt engineering, which is regarded as one of the most resource-efficient approaches to using LLMs. In particular, prompting refers to designing input prompts to guide pre-trained LLMs to generate desired outputs. It takes advantage of the inherent inference capabilities of pre-trained LLMs, and avoids the complexity of dedicated model training or fine-tuning. Compared with other approaches, prompting-based techniques have several key advantages: 1) Resource-efficient: Compared with pre-training and fine-tuning, prompt engineering techniques have much lower requirements on computational resources like a single GPU. Such a resource-efficient approach can be easily deployed in many network devices. 2) Higher flexibility: Prompting-based methods can quickly adapt to various tasks by crafting the corresponding demonstrations and queries. It indicates an efficient method to customize LLMs to address a wide range of network tasks. 3) Fast implementation: Prompt engineering relies on the inference capabilities of LLMs, and it avoids the time cost of updating LLM parameters. Therefore, the low response time can better handle low-latency services in wireless networks.
LLMs have been discussed in several existing studies, but they mainly focus on system-level discussions and module designs, e.g., edge intelligence [5], grounding and alignment[6], 6G security[7], and our previous work also investigates LLM-enabled power control and traffic prediction [8, 9]. However, this work is different from existing studies by systematically exploring prompt engineering and wireless network applications, providing detailed prompt designs and specific case studies. The main contributions include:
1) Firstly, it presents in-depth analyses of LLM fundamentals such as pre-training, fine-tuning, RAG, and prompt engineering. We compare these techniques regarding main features, computational resources and time costs, potential difficulties, and wireless network applications. Then, it overviews multiple prompting techniques, such as in-context learning, chain-of-thought, prompt-based planning, and self-refinement, and we further investigate how various prompting techniques can adapt to wireless networks in terms of main features, advantages, and potential issues.
2) We propose two novel prompting techniques, namely iterative prompting and self-refined prompting, aiming to address network optimization and prediction problems, respectively. Specifically, iterative prompting can learn from previous experience and exploration to improve LLM’s performance on target tasks iteratively, which is mainly designed for network optimization problems. By contrast, self-refined prompting allows them to correct their outputs through iterative feedback and refinement demonstration prompts, aiming to address network prediction tasks. These schemes rely on LLM’s inference capabilities, and such resource-efficient techniques align well with the limited computational capacities of network devices. In addition, our proposed algorithms can be easily generalized to various wireless network applications, and the case studies show that they achieve satisfactory performance in network power control and traffic prediction problems.
II LLM Fundamentals towards Wireless Network Applications
Figure 1: Illustration of different prompting techniques.
This section introduces various approaches for applying LLMs to wireless networks, including pre-training, fine-tuning, RAG, and prompt engineering. Specifically, Table I compares these techniques in terms of main features, advantages, potential difficulties, computational resources and time costs, and possible wireless network applications.
Firstly, pre-training is a foundational step for LLM development, which includes data collection, filtering, and model training. Pre-training enables LLMs to develop fundamental capabilities such as comprehension, reasoning, and instruction following. However, it requires considerable computational resources such as NVIDIA H100100100100-$80$G or A100100100100-$80$G for implementation, and the training process may take weeks and months. Therefore, despite the great benefits, pre-training a wireless-specific LLM from scratch can be extremely time-consuming with huge costs of computational resources.
By contrast, fine-tuning is a more affordable approach for LLM implementation. In particular, fine-tuning indicates adapting a pre-trained general-domain LLM to specific tasks by partially updating the model weights to fit smaller and domain-specific datasets such as 3GPP standards. The fine-tuning costs are related to the LLM model size and fine-tuning methods. For instance, fully fine-tuning a 7B LLM model using 8 A100 GPUs may take 100-200 GPU hours for a 100M token dataset, e.g., SPEC5G dataset with 134M words[10] and 3GPP dataset Tspec-LLM with 534M words[11]. Similarly, RAG is another promising direction for practical LLM applications for communication networks. It integrates LLMs with an external knowledge base, and then the LLM can retrieve related information from the external knowledge source during the inference. RAG can significantly improve the quality and accuracy of the generated content, since the LLM has access to the most relevant and up-to-date information.
Compared with previous techniques, prompt engineering has much lower computational costs and faster implementation. Instead of updating model parameters, it relies on LLM’s inherent capabilities, designing specific input prompts to guide the generation of pre-trained LLMs. Without the model updating burdens, prompt design and testing only need a few GPUs, and the time costs are significantly lower than fine-tuning or RAG. Due to its low costs, it can also quickly adapt to dynamic network environments, handling various network tasks with much lower response time.
The above analyses show that prompt engineering is one of the most efficient approaches to using LLMs, which aligns well with several crucial wireless network features. Firstly, many network devices are computational and energy resource-constrained, and prompt engineering has low requirements for computational resources, which can be easily deployed and implemented. Secondly, prompt engineering allows human language-based instructions, e.g., network-specific task descriptions, questions, and solutions. By integrating natural language, network intelligence will become more accessible, especially for operators without professional AI knowledge. In addition, it has higher design flexibility. With proper demonstrations, prompts can be quickly designed and applied to various network tasks, making rapid responses to network dynamics. Therefore, it is crucial to investigate the application of prompting techniques to wireless networks, showing a promising way to utilize generative AI.
TABLE II: Summary of different prompting techniques for wireless networks.
Prompting techniques Main features Advantages Potential issues Wireless network applications
In-context learning In-context learning empowers LLMs to learn from contextual information provided in task-specific instruction and demonstrations. These prompts are designed with relevant, task-specific examples that are converted into natural language comprehensible to LLM. LLMs can be generalized to perform various learning tasks by utilizing existing knowledge from pre-training data and acquiring new task-solving strategies from contextual demonstration prompts. In-context learning is a fundamental and crucial capability of LLMs. Demonstrations greatly impact in-context learning, which should be carefully designed, including the selection, format, and order of examples, and then transform the selected examples into a formatted prompt. In-context learning is a promising approach to applying LLMs to wireless networks, using previously accumulated solutions as input prompts to address new network issues.
Chain-of-thought Chain-of-thought is an advanced strategy to improve LLM’s capabilities in complex reasoning, which incorporates sequences of intermediate reasoning steps into demonstration prompts. It can guide LLMs to better understand the logical steps from questions to answers. LLMs can solve complex problems more transparently and logically through chain-of-thought demonstration prompts, such as “Let’s think step by step”. Existing studies have shown that chain-of-thought can significantly improve LLM’s performance in diverse tasks. Despite the advancement, chain-of-thought prompting still faces issues like incorrect reasoning and instability. Although LLMs generate correct final answers, their inherent instability may provide incorrect and misleading intermediate reasoning steps. The step-by-step chain-of-thought prompts can make LLMs better adapt to complicated wireless environments, which may require step-by-step reasoning, planning, and decision-making, e.g., project code generation with multi-step scheduling.
Prompt-based planning Different from in-context learning and chain-of-thought, prompt-based planning aims to address more complicated scenarios such as multi-hop question answering. It breaks complex tasks down into manageable sub-tasks. In-context learning and chain-of-thought are more suitable for single-task implementation. By contrast, LLMs can make complex tasks more manageable by using their planning capabilities, lowering the overall difficulties. Currently, LLM-based planning requires dedicated analyses to decompose a complicated task into multiple sub-tasks. Zero-shoot and automated task decomposition is still the main difficulty. Wireless networks are complicated systems and planning is a crucial capability for addressing many network tasks. LLM-enabled automated task decomposition can be extremely useful for long-term network management and project development.
Self-refinement Considering that LLMs could initially generate incorrect answers, self-refinement allows them to correct their outputs through iterative feedback and refinement demonstration prompts. Through iterative feedback generation and prediction refinement, LLMs can self-enhance their outputs with strong effectiveness and scalable capabilities. To avoid repeating previous mistakes, all historical outputs are iteratively appended as the inputs of LLMs. It may lead to increasingly lengthy demonstration prompts and unavoidable resource consumption. Self-refinement will significantly save human efforts when using LLMs to address network tasks, since LLMs can constantly improve the previous network policy automatically.
III Prompt Engineering for Wireless Networks
Figure 2: Iterative prompting schemes for wireless network optimization
This section presents several prompting techniques and network applications, i.e., in-context learning, chain-of-thought, prompt-based planning, and self-refinement methods.
Firstly, as shown in Fig.1, in-context learning is a fundamental and crucial capability of LLMs. It refers to the process of learning from formatted natural language instructions and demonstrations, and then improving the performance on target tasks. Given contextual demonstration prompts, LLMs can address various downstream tasks by using existing knowledge from pre-trained data. Demonstrations are of great importance for in-context learning, serving as critical references for LLMs to learn from. Therefore, they should be carefully selected, formatted, and designed, which may require professional understanding and knowledge. In-context learning is a promising approach to applying LLMs to wireless networks, which can take advantage of previous network solutions as demonstrations to address unseen network issues.
Meanwhile, wireless networks are complicated systems, and one task may include several logic and reasoning steps. For instance, many network optimization problems include several elements, e.g., base stations, transmission power, bandwidth, and network users. Using simple input-output pairs as demonstrations cannot provide the logical reasons behind these examples, and LLMs may have difficulty in learning. To this end, chain-of-thought is proposed as an advanced strategy to enhance LLM’s performance in complex reasoning tasks. Specifically, chain-of-thought will provide explanations for each logic step in the prompt demonstrations: “Let’s think step by step. Given the current base station transmission power $\{BS\_power\}$, using the Shannon capacity equation, the transmission data rate is $\{user\_rate\}$. Then we compare the achieved rate with the target data rate, and $\{user\_rate\}$ is lower. Therefore, we may need to increase the transmission power.” This example includes two steps, “using Shannon capacity equation” and “compare the achieved rate with the target data rate”, then the final decision is to “increase the transmission power”. With these two-step explanations, LLMs can easily capture the relationship between transmission power and target data rate, and generate better replies.
In-context learning and chain-of-thought are mainly designed for single tasks, which may have difficulty tackling more complicated tasks, e.g., collecting information from multiple sources and producing multi-step reasoning to generate a comprehensive answer. Many complex network tasks need to be broken down into more manageable sub-tasks. For example, the network configuration usually involves multiple inter-related network elements, and network project development may consist of a series of coding and testing tasks. In this case, prompt-based planning aims to generate a group of sub-tasks, which will be executed one by one. This planning capability is crucial for handling many large-scale network tasks.
In addition, LLMs may generate incorrect answers at the initial attempts, and external feedback is essential to improve its replies. Self-refinement allows LLMs to self-improve their previous outputs by generating feedback and refinement prompts. In particular, it indicates that LLMs can constantly improve their previous outputs automatically without human intervention. Self-refinement can significantly save human effort in improving LLM’s initial replies, especially considering the complexity of network tasks.
Finally, Table II compares the above techniques in terms of main features, advantages, potential issues, and wireless network applications. Note that these methods can be combined to address complex tasks comprehensively. For instance, using self-refinement in prompt-based planning to save effort in the evaluation phase, and integrating chain-of-thought to improve the multi-step logic reasoning in sub-task decomposition.
IV Iterative Prompting for Wireless Network Optimization
Optimization techniques are of great importance for wireless network management, which has been extensively studied for various network tasks, e.g., power control, resource allocation, user association, task offloading, etc. Considering its crucial importance, this section proposes a novel iterative prompting scheme for network optimization, in which we prompt LLMs iteratively to improve the performance on optimization tasks.
IV-A Natural Language-based Task Description
Firstly, as shown in Fig. 2, the target task is described by formatted natural language, including task goals, definitions, and extra rules, providing fundamental task information to the LLM. For instance, the task goal describes the problem as a “decision-making task”, and it introduces the set of control variables as “candidate decisions”. Then, it highlights the importance of “environment states”, indicating that the LLM has to consider specific environment variables for decision-making. After that, it introduces the current target task by “here is a new case to solve, and the current environment states is…”. Moreover, the final rules specify that “select from {Decision_set} based on above examples”.
With these formatted task descriptions, the LLM can capture the key elements of the decision-making task, focusing on hidden patterns between environment stats and corresponding decision variables. Such a design can also lower the understanding difficulties caused by professional network knowledge, since general domain LLMs such as GPTs and Llama are not specifically pre-trained for network applications.
Figure 3: Self-refined prompting for network traffic prediction. Here, we present an example of hourly traffic prediction for 2013-11-08 using hourly historical traffic from 2013-11-07. Due to space constraints, we show traffic only for certain hours, while traffic is available each hour during interactions with an LLM. The specific feedback related to initial predictions is incorporated into the feedback prompt. The LLM refines its predictions by following the actionable steps outlined in the refinement prompt. Due to space constraints, we only show feedback generation and prediction refinement at one iteration.
IV-B Iterative Prompting Scheme
As illustrated in Fig. 2, the above task description will become input prompts for the LLM at network edge cloud, and the LLM selects a specific Decision_n from the set of candidate decisions {Decision_set}. Then, the network management decisions will be sent to the wireless network environment for implementation, e.g., resource allocation, user association and transmission power level decisions. The network operation results be collected as a new example, including environment states, selected decisions, and rewards that are related to optimization objectives and network metrics.
After that, the accumulated examples consist of an experience pool, recording all previous explorations and examples. They will be further evaluated to select examples that are related to the target task. For instance, a straightforward strategy is to select previous examples with similar environment states as current states, e.g., similar user numbers and channel conditions. Then the previous decisions and rewards become useful references for LLM decision-making. However, note that many wireless environment states are continuous variables, e.g., channel state information or user-base station distances, and therefore it is unlikely to find examples with exactly the same states. Therefore, the evaluation and selection of examples are crucial in the proposed scheme.
Finally, the selected examples, i.e., recommended and inadvisable demonstrations, are integrated into the task description, providing useful references for LLM decision-making. Fig. 2 has presented steps 1 to 6 of the defined prompting scheme, and it can iteratively explore new decisions and accumulate new experiences. Therefore, the LLM can constantly learn from previous explorations, and then improve its decisions accordingly. Such an iterative prompting method can rapidly adapt to dynamic wireless environments. Meanwhile, the proposed scheme in Fig. 2 has several advantages over conventional machine learning approaches: 1) No model parameter updating and fine-tuning is required in this scheme, which has much lower complexity than existing algorithms such as deep reinforcement learning (DRL); 2) It utilizes natural language for network management, making network control and optimization more accessible with minimum requirements on mathematical knowledge and equations; 3) LLMs can provide reasonable explanations for their decisions, and help humans understand complicated network systems, while non-previous algorithms have such ground-breaking capabilities.
V Self-refined Prompting for Network Prediction
Prediction tasks are critical in wireless networks, using historical network data to predict future trends and behaviours such as network traffic, user demand, channel states, and device status. To showcase the potential of prompting techniques, this section presents a self-refined prompting method for network prediction problems, i.e., network traffic prediction.
V-A Prediction Task Description
A crucial step in using LLMs for prediction problems is to convert numerical values into natural language sentences. Fig. 3 considers network traffic prediction as an example, and illustrates the demonstration prompt, data prompt, and query prompt. In particular, the traffic prediction task is framed in a question-answer format, adhering to a template-based description. The demonstration prompt is first incorporated to guide the LLM in making traffic predictions by leveraging contextual task descriptions for next-day traffic prediction. Then, the LLM predicts future traffic by responding to questions in the query prompt, based on historical traffic data provided in the data prompt, i.e., previous hourly network traffic.
V-B Feedback Generation and Prediction Refinement
Initially, LLMs may generate incorrect predictions, and this subsection utilizes self-refined prompting to incorporate feedback demonstrations to provide comprehensive and valuable insights, aiming to improve the initial results [9]. As shown in Fig. 3, the feedback prompt is designed to encompass prediction performance, prediction format and completeness, and prediction method, which are shown as follows:
  • Overall performance: The same LLM evaluates the overall prediction performance, e.g., Mean Absolution Error (MAE) between ground truth and predictions.
  • Periodical performance: Considering that time-series data can be represented through sine and cosine functions in the real domain [12], the same LLM is tasked with projecting both ground truth and predicted traffic onto these functions, respectively. By matching the projection of ground truth, predictions are expected to capture the periodic nature of ground truth accurately, such as fluctuations during peak and off-peak traffic periods.
  • Format and completeness: The predictions should match the format of ground truth as well as be complete for each timestamp, e.g., 24 values for daily traffic prediction.
  • Prediction method: The prediction method is summarized in the feedback demonstration prompt, which calls for adopting more advanced and accurate methods to enhance performance further. To avoid additional computational costs, while the LLM may suggest developing a neural network as a potential improvement method, this recommendation is disregarded.
Given the above instructions, the refinement demonstration prompts aim to provide specific actionable steps associated with feedback demonstration prompts. This allows the same LLM to self-refine previous predictions by adhering to the detailed feedback outlined in feedback demonstration prompts. Note that the process of feedback generation and prediction refinement is iteratively conducted until the prediction performance converges, whereas the inference is completed without engaging in the feedback generation and prediction refinement. Additionally, to prevent the repetition of previous incorrect predictions, the history of predicted traffic, feedback, and refinement demonstration prompts are also included as inputs to the LLM at each iteration.
To summarize, the proposed prompting-based prediction method can constantly improve its prediction performance with self-refinement. Compared with conventional techniques such as long short-term memory (LSTM) networks, the prompting-based approach avoids the complexity of dedicated model training for specific datasets. It is regarded as a plug-and-predict method, providing instant prediction results without parameter fine-tuning. However, the iterative appending of all historical predictions may increases the length of demonstration prompts, potentially leading to higher computational costs and resource consumption.
(a) Optimization Case Study: improved power consumption of Llamas.
(b) Optimization Case Study: service quality comparison with training.
(c) Prediction Case Study: average MAE and MSE comparison.
(d) Prediction Case Study: predicted traffic comparison.
Figure 4: Performance comparison in network optimization and traffic prediction case studies.
VI Case Studies
VI-A Simulation Settings
This section presents two case studies on network optimization and prediction tasks: 1) Network Optimization: This case study involves a base station power control problem, which is a fundamental optimization task that has been widely investigated in wireless network fields. The considered problem formulations can be easily found in [13], which aims to minimize the base station power consumption and meanwhile maintain an average data rate threshold for users. We consider three adjacent base stations and the associated user numbers dynamically change from 5 to 15 each, the average data rate threshold is 1.5 Mbps/per user, and the channel gain applies 3GPP urban network models. ① Iterative prompting: The simulation includes 2 LLM models: Llama3-7b-instruct as a small-scale model, and Llama3-70b-instruct as a large-scale model. ② DRL baseline: DRL is included as an optimal baseline, which has been widely used to address network optimization problems.
2) Network Prediction: This case study considers a well-known network traffic dataset from the city of Milan [14]. It records the time of user interactions and the base stations managing these interactions. Each base station covers an area of approximately $1km\times 1km$, resulting in a total of 8,923 samples. ① Self-refinement: Our proposed self-refined prompting utilizes GPT-4 as the foundation model for one-day-ahead traffic prediction using historical traffic from the previous day. ② LLM-related baselines: Meanwhile, we include GPT-4 and GPT-3.5 without self-refinement as two LLM-related baselines, better demonstrating the capabilities of self-refinement. ③ Conventional baselines: We also consider autoregressive integrated moving average (ARIMA) and LSTM as two conventional baselines. We deploy LSTM as an optimal baseline, since it has been specifically trained on the target network traffic dataset. By contrast, ① Self-refinemen and ② LLM-related baselines have no previous knowledge on the target task.
VI-B Simulation Results and Analyses
1) Network Optimization Case Study: Fig. 4(a) and 4(b) present the optimization task results. Specifically, one can observe that the average power consumption is constantly improved with the increasing number of episodes in Fig. 4(a). It indicates that the proposed scheme in Section IV enables LLMs to learn from previous examples and experience, and then improve its performance on target tasks iteratively. Fig. 4(a) also shows that LLMs present comparable performance as the DRL algorithm, which is achieved without model training or fine-tuning as in conventional ML algorithms. Meanwhile, Fig. 4(a) compares the service quality of Llama3-7b, Llama3-70b, and DRL, which means the probability of violating the preset average data rate threshold of 1.5 Mbps/per user. It highlights that LLMs can maintain high service quality at the beginning of the training process, while the DRL algorithm has a lower service quality in the initial exploration phase. Fig. 4(a) reveals the potential advantage of LLMs in terms of learning efficiency. Given the dedicated designed task description and examples, LLM can quickly learn from human instructions and previous demonstrations. By contrast, the DRL approach relies on exploration and value updating to understand the task, and the low sampling efficiency is a well-known issue of DRL methods.
2) Network Prediction Case Study: Fig. 4(c) compares the average MAE and MSE metrics of various techniques. It is observed that the self-refined prompting method significantly outperforms ARIMA, GPT-3.5, and GPT-4, demonstrating the capabilities of our proposed self-refined prediction method. The self-refined prompting method exhibits notable improvements over GPT-3.5 and GPT-4, with MAE reductions of 29.72% and 17.09%, which underscores more powerful generalization capability on unseen datasets through iterative feedback generation and prediction refinement processes. GPT-3.5 shows a worse performance due to its outdated architecture and designs, which have been observed in many existing studies. In addition, LSTM shows the best performance as an optimal baseline. However, it is worth noting that LSTM has been specifically trained on the target dataset, while all LLM-based methods are zero-shot predictors. Furthermore, similar results can be observed in Fig. 4(d), which illustrates the one-day ground truth traffic and the predicted traffic of various methods. The self-refined prediction method more closely matches the ground truth traffic, indicating strong scalability of predictions across various base stations. By contrast, GPT-4 shows an obvious mismatch from 9:00 to 15:00.
In summary, the simulation results in Fig. 4 demonstrate that LLMs have great potential in handling network management, i.e., optimization and prediction tasks. They present satisfied performance as conventional machine learning algorithms such as DRL and LSTM, while avoiding extra model training and fine-tuning complexity. LLMs also show higher learning efficiency than existing techniques, and they allow human language-based input, showing the possibility of natural language-based future network management.
VII Conclusion
The progress of generative AI and LLMs brings promising opportunities for next-generation wireless networks. This work provides a comprehensive overview of prompt engineering for network applications. It presents in-depth analyses of different prompting techniques such as in-context learning, chain-of-thought, and self-refinement. In addition, this work proposes two novel prompting schemes for network optimization and prediction problems, and the case studies show that the proposed methods can achieve satisfactory performance.